40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11219     Entry time: Wed Apr 15 02:26:41 2015
Author: ericq 
Type: Update 
Category: LSC 
Subject: Attempted DRMI + ALS arms 

I spent some time tonight trying to revive DRMI locking, with the arms held off on ALS. Not much news, I haven't been able to get more than a few short spurts of resonance, using 1F signals.

I did use SRY to measure the BS->SRCL coupling by exciting each mirror and looking at their relative coupling to AS55Q. I found that we should use a value of +0.28 +-0.01 in the MICH->SRM element.

ELOG V3.1.3-